### v2

# Flux-Activated Resonant Control of a Bosonic Quantum Memory

Abstract
Universal control of bosonic degrees of freedom provides a hardware-efficient route for quantum information processing with high-dimensional systems. Bosonic circuit quantum electrodynamics (cQED), which leverages auxillary transmons to coherently control long-lived superconducting cavities, is well suited to this goal. However, such systems are traditionally operated in the dispersive regime, where the nearly degenerate cavity transitions prohibit the direct addressability of individual excitation levels of the bosonic mode and increase gate complexity. Here, we achieve direct oscillator control by dynamically accessing the resonant Jaynes–Cummings (JC) interactions, implemented with a hardware that integrates on-chip broadband magnetic flux delivery with a bosonic memory housed in a 3D superconducting cavity with lifetime exceeding 0.5 ms. We demonstrate deterministic preparation of Fock states and their superpositions within 100s of nanoseconds by directly climbing the JC ladder, and realise efficient arbitrary rotations between any pair of Fock states. Our resonant control scheme provides an analytical method for manipulating the entire Hilbert space of the bosonic mode at rates fundamentally more favourable than traditional dispersive strategies. This on-demand access to JC interactions opens a promising path toward realising robust Fock-basis qudits and harnessing the rich dynamics of high-dimensional bosonic systems for quantum information processing.
Effective control over a large Hilbert space is essential for both practical quantum information processing [61, 76, 22] and explorations of the rich underlying concepts of quantum physics [2, 21, 1]. A promising route towards this goal is to leverage the intrinsically large Hilbert space of quantum harmonic oscillators. Notably, bosonic circuit QED systems use 3D superconducting cavities coupled to auxillary transmons to realise highly controllable and coherent oscillators. Such devices have demonstrated Fock state preparation up to 100 photons [11] and cat states with an average photon population of $1024$ [52]. Compared to parallel pursuits using discrete $d$-level systems (qudits) [37, 71], such as molecular spins [58, 54], optical photons [72, 28], trapped ions [42, 60], silicon nuclear spins [75] and superconducting artificial atoms [73, 41, 9], encoding information in higher levels of the oscillator does not introduce additional decoherence channels and is readily compatible with a broad range of existing bosonic quantum error correction schemes [65, 16].
In cQED systems, universal control of the bosonic mode is typically achieved via dispersive coupling to a nonlinear auxillary element, commonly in the form of a transmon qubit [5]. In this regime, oscillator transitions are close in frequency and not individually addressable, as illustrated (Fig. 1a). Therefore, arbitrary operations on the oscillator state often require driving the transmon and the cavity simultaneously with complex numerically optimised waveforms [32, 18, 17]. However, these numerical pulses are susceptible to parasitic dynamics, and their ‘black-box’ nature often complicates effective troubleshooting. Although operations can also be implemented using higher-order parametric processes without numerical optimization [23, 29], they are limited by the achievable coupling strength or drive-induced errors [69, 45]. Therefore, an analytical and versatile control method that directly addresses individual cavity transitions would fill an important gap for bosonic quantum information processing.
In this work, we use flux-activated resonant Jaynes-Cummings (JC) interactions to access number-selective cavity transitions and implement arbitrary yet analytically simple control of a bosonic memory mode. These dynamics are enabled by adjusting the transmon-cavity detuning on demand via a broadband on-chip flux line. This allows us to rapidly switch the coupled system between a highly nonlinear resonant dynamics with well-resolved JC transitions and the dispersive regime with suppressed cavity photon loss. We protect the coherence of the bosonic memory against flux line leakage by incorporating on-chip low-pass filtering over the frequency band occupied by the cavity and the transmon. As a result, the lifetime of the bosonic mode reaches $T_{1}=590\pm 40\,\mu\textnormal{s}$ ($Q\sim 2.6\times 10^{7}$) when the transmon is tuned to be dispersively coupled – more than a two-fold improvement over previous implementations using the magnetic hose [68, 3]. We leverage these hardware capabilities to demonstrate the direct creation of Fock states and their superpositions by sequentially climbing the JC ladder. Furthermore, we demonstrate state transfer between any two chosen energy levels of the cavity using multi-tone pulses and realise arbitrary rotations between them. This resonant control scheme can readily be applied to the general class of spin-boson systems, such as opto-acoustic modes coupled to artificial atoms [4] or motional degrees of freedom of ions controlled by a spin [48]. Thus, our results establish a valuable building block for efficient universal control and quantum information processing using bosonic modes in different physical platforms.
Figure 1: Comparison of the dispersive and resonant regimes. (a) The bosonic transitions in the dispersive regime are degenerate within each qubit manifold and are separated by dispersive coupling strength, $\chi$, between the $|\mathord{\downarrow}\rangle$ and $|\mathord{\uparrow}\rangle$ manifolds. (b) This degeneracy is lifted in the resonant JC regime, with level splittings $\propto g\gg\chi$, enabling fast selective excitation of individual bosonic levels. (c) The frequency spectra of both coupling regimes shown on the same axis.
The resonant control technique explored in this work is based on the theoretical proposal by Strauch [63], which adiabatically connects the eigenstates of the resonant JC Hamiltonian to the product eigenbasis of the decoupled circuits. It improves upon earlier works [38, 20] that rely on sequentially turning on and off the spin-boson coupling.
The dynamics of a spin-boson system on resonance is well approximated within the rotating wave approximation by the JC Hamiltonian [26]
$\hat{H}_{\textnormal{JC}}=g\left(\hat{a}^{\dagger}\hat{\sigma}^{-}+\hat{a}\hat{\sigma}^{+}\right),$ (1)
where $\hat{a}$ ($\hat{a}^{\dagger}$) is the annihilation (creation) operator of the boson in the form of a superconducting cavity, $\hat{\sigma}^{-}$ ($\hat{\sigma}^{+}$) is the lowering (raising) Pauli operator of the spin, realised with a transmon qubit, and $g$ is the coupling strength between them. The $\hat{H}_{\textnormal{JC}}$ eigenstates are the ground state $|0,\,\downarrow\rangle$ and the $N$th-excited levels
$|N\pm\rangle=\frac{1}{\sqrt{2}}\left(|N,\,\downarrow\rangle\pm|N\mathrm{-}1,\,\uparrow\rangle\right),$ (2)
where $|N,s\rangle$ is the product of the cavity Fock state $|N\rangle$ and the transmon state $|s=\,\downarrow$$\uparrow\rangle,$ with eigenfrequencies $\omega_{|N\pm\rangle}=\pm\sqrt{N}g$. This photon-number dependence of $\omega_{|N\pm\rangle}$ leads to the direct addressability of the transitions as illustrated in Fig. 1(b).
We are particularly interested in the red and blue JC sidebands
$|N+\rangle\,\overset{\omega_{N,r}}{\longleftrightarrow}\,|(N\mathrm{+}1)-\rangle,$ (3)
$|N-\rangle\,\overset{\omega_{N,b}}{\longleftrightarrow}\,|(N\mathrm{+}1)+\rangle,$
with transition frequencies $\omega_{N,r}=-(\sqrt{N\mathrm{+}1}+\sqrt{N})g$ and $\omega_{N,b}=(\sqrt{N\mathrm{+}1}+\sqrt{N})g$ relative to the rotating frame, respectively. Transitions $\omega_{1,b}$, $\omega_{2,r}$, and $\omega_{3,b}$ are shown with orange, green, and teal, in Fig. 1. This set of transitions, together with ground state excitations $|0,\,\downarrow\rangle\leftrightarrow|1\pm\rangle$, can selectively connect any two eigenstates of the resonant JC Hamiltonian with a spectral resolution on the order of $g$ (Fig. 1c). That makes the timescale of our protocol fundamentally more favourable compared to control schemes using the dispersive coupling, as discussed in [64].
To translate this fast control on the JC levels to effective operations on the bosonic mode alone, we simply need to adiabatically decrease the transmon frequency away from the cavity resonance. This transforms the coupled eigenstates in Eq. (2) into products of cavity and transmon eigenstates: $|N+\rangle\to|N,\,\downarrow\rangle$ and $|N-\rangle\to|N\mathrm{-}1,\,\uparrow\rangle$. The red sidebands in Eq. (3) map to the transitions $|N,\,\downarrow\rangle\leftrightarrow|N,\,\uparrow\rangle$, similar to the standard transmon excitations in the dispersive regime [59]. In contrast, the blue sideband, $|N\mathrm{-}1,\,\uparrow\rangle\leftrightarrow|N\mathrm{+}1,\,\downarrow\rangle$, is only directly accessible in the resonant JC dynamics [63]. Therefore, the on-demand access to the resonant JC dynamics, together with the ability to adiabatically detune the transmon from the oscillator, allows us to orchestrate any desired operations on bosonic mode.
To leverage this resonant control scheme, we construct a new bosonic circuit QED hardware that can simultaneously provide strong and broadband flux control while protecting the cavity from decay (Fig. 2a). Our device features a superconducting stub cavity as the long-lived quantum memory coupled to a transmon qubit with a coupling strength $g/2\pi=12.2\pm 0.1\,\textnormal{MHz}$. It is combined with a lithographically defined on-chip flux line placed in close proximity to the SQUID loop with area $470\,\mu\textnormal{m}^{2}$. Flux delivery is highly efficient, featuring a mutual inductance of $0.387\,\Phi_{0}/\textnormal{mA}$. This allows us to minimise the loop area and hence reduce the impact of flux-noise on the transmon coherence times [64]. In this particular device, we measure transmon $T_{1}\approx 23\,\mu\textnormal{s}$ and $T_{2}^{\textnormal{echo}}\approx 3.5\,\mu\textnormal{s}$ in the lower sweet spot. The flux line bandwidth reaches at least $1.3\,\textnormal{GHz}$, allowing fast transmon frequency adjustment in the nanosecond timescale. Although we did not calibrate at higher frequencies due to the presence of a 2 GHz low-pass filter on the coaxial line, our finite-element simulations predict the bandwidth $\sim\textnormal{4}$ GHz, limited by the on-chip low-pass filter [64].
Figure 2: Architecture for robust flux-delivery in a high-Q bosonic circuit QED device. (a) High-Q $\lambda/4$ coaxial cavity coupled to an asymmetric SQUID-transmon (shown in bottom-left inset). The flux bias is delivered by a transmission line wirebonded to an SMA port at the opposite end of the chip. The flux line and a readout resonator partially adopt a coplanar waveguide (CPW) architecture. They are equipped with two 33rd-order Chebyshev low-pass filters in series (schematics shown in top-right inset) and a Purcell filter, respectively, to protect the cavity and the transmon against decay. (b) The cavity $T_{1}$ is measured while the transmon is parked at a few selected flux points with different values of $\Delta=(\omega_{\textnormal{cav}}-\omega_{\textnormal{T}})/2\pi$, showing an increase in $T_{1}$ as the transmon-induced decay is suppressed at larger detunings.
This CPW-based chip design is carefully optimised to reduce the energy loss of cavity and transmon through the flux line. In comparison to the microstrip flux line shown in a recent work [40], our strategy provides a more compact solution that allows convenient integration of additional circuits such as the readout resonator and cascaded filtering on the same chip. We use two 33rd-order low-pass Chebyshev filters [56] with cutoff frequency of 4 GHz in series, schematically depicted in top-right inset of Fig. 1a. Based on finite element simulations, this structure achieves $98\,\textnormal{dB}$ attenuation at the cavity frequency of $\omega_{\textnormal{cav}}/2\pi=6.868\,\textnormal{GHz}$ and at least $50\,\textnormal{dB}$ attenuation in the transmon frequency range of $\omega_{\textnormal{T}}/2\pi=5.894-7.634\,\textnormal{GHz}$. As a result, energy loss through the flux line is heavily suppressed for both quantum memory and transmon. We verify this by preparing the cavity in a coherent state and tracking the vacuum state $|0\rangle$ population over time at a few different cavity-transmon detunings as shown in Fig. 2b. The lifetime reaches a maximum of $T_{1}=590\pm 40\,\mu\textnormal{s}$ at a detuning of $\Delta=(\omega_{\textnormal{cav}}-\omega_{\textnormal{T}})/2\pi=260\,\textnormal{MHz}$, at which point the inverse-Purcell decay is suppressed. These photon loss values are comparable to those of typical 3D bosonic superconducting devices without flux integration.
Our device combines a highly coherent bosonic mode with fast tunability of a nonlinear auxiliary element, making it an ideal testbed for resonant control protocols. We first showcase the efficacy of this control framework by preparing arbitrary superpositions of Fock states with a set of analytical pulses. For example, we prepare $\frac{1}{\sqrt{2}}\left(|0\rangle|+|3\rangle\right)$ (Fig. 3a) using a pulse sequence consisting of a $\pi/2$-pulse from $|0\rangle$ to $|1+\rangle$ followed by two $\pi$-pulses to $|2$-$\rangle$ and $|3+\rangle$. The frequencies of each transition are determined using standard spectroscopy experiments, and the pulses durations are chosen to maximise the photon-number parity contrast after climbing each step of the JC ladder [64]. We calibrate the optimal speed of the adiabatic detuning by first preparing $|1+\rangle$ on resonance, followed by a flux pulse with a smooth, variable-duration ramp until the system is in a strong dispersive regime with $\Delta=100\,\textnormal{MHz}$. We then directly measure the cavity $|1\rangle$ population using the photon-number-resolved spectroscopy and choose a minimal ramp duration of $200\,\textnormal{ns}$ that maximises the population of $|1\rangle$. We further showcase the ease of orchestrating arbitrary multi-component superpositions of Fock states by creating $\tfrac{1}{2}|0\rangle+\tfrac{\sqrt{3}}{\sqrt{8}}\left(i|2\rangle+|4\rangle\right)$ (Fig. 3b) using the same method. The phase of the superposition is implemented by simply changing the phase of the pulse addressing the corresponding transition. This technique can be directly used to create arbitrary logical states superpositions in the binomial encoding [50, 36]. The total pulse sequence to prepare the state in Fig. 3a (b) takes 640 ns (688 ns), and the corresponding fidelity of the final state is estimated as $\textnormal{93}\pm\textnormal{3}\%$ ($\textnormal{89}\pm\textnormal{2}\%$) from density matrix reconstruction [34, 35], limited by transmon decoherence [64].
As an alternative to sequentially climbing the JC ladder, several transitions can also be activated simultaneously to realise more complex dynamics. For instance, the analog of the spin angular momentum $J_{x}$ over a subspace of dimension $d$ of the JC Hamiltonian can be constructed by driving each transition with amplitude
$A_{n}=2\eta_{n}A\sqrt{n\left(d-n\right)},\,\,\,1\leq n<d,$ (4)
where A is the overall pulse amplitude and $\eta_{n}$ is an experimentally calibrated scaling factor of the drive strength associated to each transition [64]. Driving $J_{x}$ for a total time of $\pi/A$ enacts pairwise state transfers between target levels. Here we use this multi-tone drive to create Fock state $|5\rangle$ from the ground state. In principle, the optimal amplitudes of each spectral component can be analytically derived to implement a perfect state transfer [63]. In practice, to improve the fidelity beyond the analytical solution, we allow $A_{n}$ to assume complex values and obtain the amplitudes using a gradient descent algorithm which maximises the state preparation fidelity for a fixed time [63, 64]. We present, in Fig. 3c, the measured Wigner function of a Fock $|5\rangle$ prepared with this method using a $800\,\textnormal{ns}$ pulse. We obtain a state fidelity of $\textnormal{75}\pm\textnormal{2}\%$, limited by both the transmon decoherence and imperfect calibrations of $\eta_{n}$ [64].
Figure 3: State preparation in the resonant regime. Left panel: Normalised Wigner plots and 1D cuts along $\mathrm{Re}[\beta]=0$ of states created using the flux-activated resonant interactions. Right panel: corresponding transitions activated to climb the JC ladder. (a) State $\tfrac{1}{\sqrt{2}}\left(|0\rangle+|3\rangle\right)$, and (b) state $\tfrac{1}{2}|0\rangle+\tfrac{\sqrt{3}}{\sqrt{8}}\left(i|2\rangle+|4\rangle\right)$ are created using sequential drives. (c) Fock state $|5\rangle$ created from vacuum by simultaneously driving multiple JC transitions. The progress bars illustrate the steps of state preparation, and the circles indicate the resulting Fock populations.
Figure 4: Givens rotation $\hat{\mathcal{R}}_{1,4}(\theta)$. (a) The $|1+\rangle$ state population is swapped with $|3$-$\rangle$ state using a $J_{x}$ pulse, which is then followed by a sideband pulse to $|4+\rangle$ with an effective rotation angle $\theta$ and a second $J_{x}$ pulse that reverts the first swap. This sequence effectively implements a gate between states $|1+\rangle$ and $|4+\rangle$ that translates into a Givens rotation between Fock states $|1\rangle$ and $|4\rangle$ after adiabatic detuning. (b) Normalised final state parity measured for arbitrary rotation angle $\theta$, showing that the protocol can create arbitrary superpositions of two Fock states. (c) Normalised Wigner function plots and Fock states populations of the cavity state for $\theta=0,\pi/4,\pi/2$, and $\pi$. The population distributions are obtained from the reconstructed density matrices and consistent with both the ideal state (target) and the outcomes simulated with real experimental parameters (black).
Furthermore, we use the multi-tone control to apply arbitrary rotations between two Fock states. This enacts Givens rotations on a qudit in Fock encoding: $\hat{\mathcal{R}}_{n_{A},n_{B}}(\theta)=\exp[-\imath\theta/2(|n_{A}\rangle\langle n_{B}|+|n_{B}\rangle\langle n_{A}|)]$. Any arbitrary operation on a qudit can be constructed via a combination of $\hat{\mathcal{R}}_{n_{A},n_{B}}(\theta)$ [63, 8], promising a simpler and faster control compared to the previous work [53]. Our implementation consists of resonantly transferring the population of a level $|n_{A}+\rangle$ to $|\left(n_{B}\mathrm{-}1\right)$-$\rangle$, and then driving a transition from the latter state to $|n_{B}+\rangle$ with an effective rotation angle $\theta$ (Fig. 4a). The sequence is concluded with a second $J_{x}$ applied in reverse, which swaps $|\left(n_{B}\mathrm{-}1\right)$-$\rangle$ and $|n_{A}+\rangle$ populations while reverting spectator levels to their original configuration. After the transmon is adiabatically detuned, this operation is equivalent to a Givens rotation between the cavity levels $|n_{A}\rangle$ and $|n_{B}\rangle$ [8]. This operation can be naturally extended to perform gates in any chosen subspace, such as the logical states of the binomial encoding, as we demonstrate in [64]. As a concrete example, we apply Givens rotations $\hat{\mathcal{R}}_{1,4}(\theta)$ between Fock states $|1\rangle$ and $|4\rangle$. The parity of the final state is measured as a function of $\theta$, showing a continuous evolution between a minimum value corresponding to $|1\rangle$ and a maximum corresponding to $|4\rangle$ (Fig. 4b). For $\theta=0,\pi/4,\pi/2$ and $\pi$, we measure the full Wigner function of the cavity as shown in the plots in Fig. 4c. The density matrix reconstruction of the Wigner measurements indicates state fidelities of $\textnormal{80}\pm\textnormal{2}\%$, $\textnormal{76}\pm\textnormal{2}\%$, $\textnormal{71}\pm\textnormal{2}\%$, $\textnormal{68}\pm\textnormal{2}\%$, respectively. This is consistent with the numerical simulation with real experimental parameters [64].
The performance of the Givens rotation and state preparation are limited mainly by transmon dephasing, followed by coherent errors and transmon $T_{1}$ [64]. The effect of each decoherence channel can be understood in terms of the matrix element of JC levels with the corresponding jump operators. Interestingly, transmon pure dephasing does not lead to dephasing of the JC levels, as the corresponding jump operator $\hat{\sigma}_{z}$ has zero matrix element $\langle N\pm|\sigma_{z}|N\pm\rangle=0$. Instead, it translates to depolarisation within each excitation manifold with corresponding matrix element $\langle N\pm|\sigma_{z}|N\mp\rangle=1$.
In the limit of long coherence times of the memory mode, the energy decay of the resonant system is predominantly defined by the transmon relaxation rate corresponding to the jump operator $\hat{\sigma}^{-}$. The matrix elements defining the loss of energy are $\langle(N\mathrm{-}1)\pm|\sigma^{-}|N\pm\rangle=\langle(N\mathrm{-}1)\mp|\sigma^{-}|N\pm\rangle=0.5$, indicating that the energy decay rate of the composite system is half that of the transmon regardless of $N$. In this implementation, the durations of the individual drives are limited by the effective lifetime of the system on resonance. This constraint results in coherent errors caused by leakages to spectator JC levels. A full master equation simulation with these error sources taken into account consistently reproduces experimental observations [64].
To enhance the overall fidelity of our resonant control scheme, we can readily adopt the known strategies to mitigate the decoherence of flux-tunable transmons, which is the dominant source of imperfection in this implemenation. For example, the transmon $T_{2}$ can be increased by reducing the flux sensitivity of the SQUID either by decreasing its area, increasing the asymmetry between the shunting Josephson junctions [24, 64], or using geometries that are inherently more noise-resilient [7]. At the same time, flux modulation can be used to engineer AC sweet spots at the cavity frequency to reduce the direct decoherence limit during the operations. Both strategies are supported by the large bandwidth and strong flux delivery of the on-chip line developed in this work. Besides greater $T_{2}$, increasing the transmon-cavity coupling strength $g$ leads to more selective sideband transitions and faster adiabatic detuning, both of which can reduce the duration of the protocol and mitigate errors. Thus, the performance of the resonant operations shown in this work is not fundamentally limited by the control mechanisms or our device architecture, and can be directly improved by further optimisation of device parameters.
In summary, our work demonstrates the control of a long-lived bosonic quantum memory through number-selective transitions in the JC Hamiltonian. This new capability is enabled by a robust flux-tunable architecture that can effectively switch the system between the dispersive and the resonant regimes. We prepare arbitrary superpositions of Fock states via either sequential or multi-tone sideband driving. Moreover, we use the JC transitions to efficiently implement arbitrary Givens rotations, providing a path for the universal control of qudits encoded in oscillators. Finally, thanks to the direct access to the JC levels, our protocol is intuitive and calibration-friendly. Moreover, it leverages a frequency selectivity that scales with the coupling factor $g$, allowing potentially faster control compared to typical dispersive operations and other parametrically activated techniques. Therefore, our work marks an important step towards more efficient control capabilities of long-lived microwave quantum memories in cQED and opens up new possibilities for quantum information processing with bosonic modes.
Acknowledgements We acknowledge the funding support from the Singapore Ministry of Education (T2EP50222-0017) and The University of Sydney - National University of Singapore 2026 Ignition Grants. F.V., A.D., N.H., acknowledge the support of the Singapore National Quantum Scholarship Scheme (NQSS). We thank Mr. Kyle Chu and Mr. Jean-Samuel Tettekpoe for their technical inputs during the project development. We thank Dr. Yao Lu and Dr. Mustafa Bakr for the valuable discussions during the writing of this manuscript.
Supplementary Materials
Flux-Activated Resonant Control of a Bosonic Quantum Memory
I Device Design
The standard bosonic cQED architecture consists of a high-Q bosonic mode hosted in a 3D superconducting cavity with convenient control and measurement tools provided by an on-chip ancillary transmon and a planar low-Q readout resonator. Our device extends this design by integrating fast-flux tunability without compromising long cavity lifetimes.
These flux control features are achieved by combining coplanar waveguide (CPW) and microstrip line architectures. We show the details of the circuit in Fig. S1. The tunable transmon is designed in a microstrip geometry, allowing it to have a large dipole moment that couples strongly to the cavity. The SQUID loop of the transmon is inductively coupled to an on-chip coil biased by a CPW flux line. The flux line is designed with an integrated microwave low-pass filter that mitigates cavity and transmon photon loss.
In this section, we explain the technical specifications and design considerations of the tunable transmon and the flux line control.
Figure S1: Chip schematic. It features ancilla DC-SQUID transmon, low-Q readout resonator with Purcell filter, and on-chip flux line with microwave low-pass filter
I.1 Tunable transmon with asymmetric SQUID
Standard transmon [31] qubits feature nonlinear inductance implemented as a Josephson junction, leading to anharmonic oscillator spectrum where one can isolate two-dimensional computational subspace with the transition frequency $\omega_{\textnormal{T}}$ given by
$\hbar\omega_{\textnormal{T}}=\sqrt{8E_{C}E_{J}}-E_{C},$ (S1)
where $E_{J}$ and $E_{C}$ stand for Josephson and capacitive energies, respectively.
In order to make such a qubit tunable, single Josephson junction should be replaced by a loop with two junctions in parallel, forming a DC-SQUID. Magnetic flux $\Phi_{e}$ threaded through the SQUID loop effectively changes the value of the inductance and, therefore, the qubit frequency, according to
$\hbar\omega_{\textnormal{T}}=\sqrt{8E_{C}\cdot 2E_{J}|\cos{\varphi_{e}}|}-E_{C},$ (S2)
where $E_{J}$ is the Josephson energy of a single junction, $\varphi_{e}=\pi\frac{\Phi_{e}}{\Phi_{0}}$ the normalised flux , and $\Phi_{0}$ the magnetic flux quantum. Typical spectrum of $\omega_{\textnormal{T}}$ as a function of the applied magnetic flux is shown in Fig. S2(a).
Figure S2: Reduced noise sensitivity of asymmetric SQUID transmons. (a) Frequency response for variable junction asymmetry factors $\gamma$. (b) Calculated transmon frequency sensitivity to flux noise as a function of flux bias.
In such tunable devices, noise in the DC bias or in the magnetic environment leads to the fluctuations of the magnetic flux. This, in turn, varies the transmon frequency and leads to significant limitations on the $T_{2}$ lifetime. Thus, it is beneficial to operate at the sweetspots, where the qubit is first-order insensitive to the flux noise. At other bias points, the qubit suffers rapidly increasing dephasing.
This sensitivity to the flux noise can be quantified as $\frac{\partial\omega_{\textnormal{T}}}{\partial\Phi_{e}}$, i.e. how much the qubit frequency reacts to the fluctuations of the magnetic flux due to the noise. From Eq. (S2),
$\frac{\partial\omega_{\textnormal{T}}}{\partial\Phi_{e}}=-\frac{\pi^{2}}{\Phi_{0}}\sqrt{2\frac{E_{C}\cdot 2E_{J}}{h^{2}}}\frac{\sin{2\varphi_{e}}}{\cos^{\frac{3}{2}}{\varphi_{e}}}.$ (S3)
The sensitivity to the flux noise grows rapidly outside of the sweetspots due to the strong dependence of frequency on the applied flux bias. Here, we intentionally suppress this dependency by making the two junctions asymmetric, i.e. having different Josephson energies [25]. The extent of asymmetry is expressed as
$\gamma=\frac{E_{J2}}{E_{J1}},$ (S4)
where $E_{J1}$ and $E_{J2}$ are Josephson energies of the junctions and $\gamma>1$ without loss of generality.
Naturally, the spectrum of the asymmetric transmon will also depend on the parameter $\gamma$:
$\hbar\omega_{\textnormal{T}}=\sqrt{8E_{C}\cdot E_{J1}\sqrt{\gamma^{2}+2\gamma\cos{2\varphi_{e}}+1}}-E_{C}.$ (S5)
We calculate the flux noise sensitivity analogously to Eq. (S3) using Eq. (S5) and obtain
$\frac{\partial\omega_{\textnormal{T}}}{\partial\Phi_{e}}=-\frac{\pi^{2}}{\Phi_{0}}\sqrt{\frac{2E_{C}E_{J1}}{h^{2}}}\frac{4\gamma\sin{2\varphi_{e}}}{\big(\gamma^{2}+2\gamma\cos{2\varphi_{e}}+1\big)^{\frac{3}{4}}}.$ (S6)
We present the simulated flux sensitivies obtained using Eq. (S3) and Eq. (S6) in Fig. S2(b). One can observe that more asymmetric transmons are less sensitive to the flux noise at the expense of a smaller frequency range. Therefore, for this particular device, we designed the SQUID to be asymmetric with a parameter $\gamma=4.02$. The resulting frequency range is 5.894 - 7.634 GHz, as shown in Fig. S3(a).
It is worth mentioning that Eq. (S2) and further analysis, strictly speaking, is only valid for the applied biasing DC current. In case of an AC flux drive, a more rigorous analysis [74, 44] must be applied to ensure an effective differential drive on the SQUID. The use of an asymmetric SQUID makes this optimisation more complicated. Although it is outside of the scope of this work, we are actively exploring alternative strategies that can balance the robustness to flux noise with the effective suppression of the common mode drive for future works that require clean AC modulation of the qubit frequency.
I.2 Flux line & on-chip microwave low-pass filter
The integration of on-demand adjustment of the transmon frequency without reducing the cavity coherence times had been an outstanding challenge for the bosonic cQED community. This feature is crucial as it enables powerful quantum information processing protocols [63, 66], as well as new opportunities for quantum simulation [27, 70] and quantum machine learning [19, 67]. However, initial attempts of the implementation of the broadband on-chip flux line encountered an issue of compromising cavity coherence [57]. In recent efforts, alternative ways of the flux delivery to the SQUID loop of the tunable ancilla have been explored  [15, 68, 3]. However, these implementations tend to be difficult to manufacture reproducibly and offer limited bandwidth ($\sim$ 100 MHz).
In this work, we demonstrate the device that combines an on-chip flux line with both effective DC and AC flux delivery with a long-lived cavity. That is achieved by implementing heavy filtering of the flux line: we incorporate compact, third order Chebyshev low-pass filter [56] with cutoff frequency of 4 GHz on the same chip. To ensure effective filtering performance and protection of the cavity mode, we concatenate two such filters, with the stub $Z_{2}$ shared between the two, as depicted schematically in Fig. S1. We leverage the CPW architecture, which allows to reach necessary impedances in much more compact design compared to the microstrip analogue [40] and to fit both filtering stages on the 4.1 mm-wide chip. Moreover, our CPW architecture also suppresses cross-talk between the flux and readout modes hosted on the same chip.
The impedance values of the series inductances and stub capacitances used in our filter are listed in Tab. I, along with the corresponding CPW line dimensions.
[…]
$Z_{0}$ 50 11 5
$Z_{1}$ 130 5 100
$Z_{2}$ 81 5 13
$Z_{3}$ 46 15 5
Table I: Impedances of the 3-rd order Chebyshev low-pass filter. Filter diagram is shown in Fig. 2 (a) of the main text.
The simulated response of the on-chip low-pass filter is shown in Fig. S3 (b). The attenuation at the cavity frequency (6.868 GHz) is 98 dB. For the entire transmon frequency band, the attenuation is at least 50 dB.
Figure S3: Filtering and spectrum. (a) A spectrum of the transmon with asymmetric SQUID. Cavity and readout resonators frequency are shown. (b) $S_{21}$ parameter of the cascaded low-pass 3-rd order Chebyshev filter.
[…]
Figure S4: Optical images of the device after electron beam lithography. (a) SQUID loop (yellow) with asymmetric junctions (white). (b) Flux line coil, biasing the transmon.
Furthermore, we incorporate a long constriction in the volume where the chip is housed, which we dub the "attenuation region". It considerably increases the cutoff frequency of the waveguide structure, enhancing the suppression of the cavity field traveling outside the ground plane. Hence, the cavity field can only propagate to the drive ports through the CPW mode, which is heavily filtered.
I.3 Readout and Purcell filter
Another advantage of the compactness of the CPW architecture, realised with the hybrid chip design, is its simplicity in integration of the readout resonator and the associated bandpass Purcell filter on the same chip. Our Purcell filter, which is shown in Fig. S1, is strongly coupled capacitively to both the readout resonator and the feed line. With this design, our readout resonator, which we place at 7.51 GHz, has coupling quality factor of $Q\approx$ 700 without introducing any notable limitations on the transmon $Q$ factor.
II Device Fabrication
The device used in this study has considerably larger footprint compared to typical transmon chips used in bosonic cQED devices due to the ground plane and the flux line with low-pass microwave filter. For practical purposes, we only use EBL for patterning the transmon, while photolithography was used for all the other parts of the chip, as summarised in the Tab. II.
# of lithography step Exposure method Material Written feature
1 EBL Nb Alignment markers
2 Photolithography Al Readout line, flux line, ground plane
3 EBL Al Transmon with the SQUID loop
Table II: Summary of the main fabrication steps. A breakdown of the processes used in fabricating respective on-chip features.
An optical microscope image of the SQUID right after EBL exposure is shown in Fig. S4(a).
Step 1: alignment markers
[…]
To clean leftover resist before the metal deposition, we apply etching for 60 seconds in Ar ($85\,\%$) – O2 ($15\,\%$) environment and evaporate 115 nm of Nb using Angstrom Engineering double-angle evaporation system. We perform lift-off via soaking the wafer in NMP at 90C for 3 hours and then sonicating in NMP, acetone and methanol for 3 minutes each.
Step 2: readout line, flux line, ground plane
[…]
Step 3: transmon with SQUID
Finally, we perform standard electron beam lithography to make the trasmon qubit. The junctions are fabricated using the bridge-free technique, presented in [39]. We then perform double-angle evaporation by first etching the resist at $\pm$20 angles for 15 seconds each. Subsequently, we deposit two layers of Al with thicknesses of 20 nm and 40 nm, separated by oxidation for 20 minutes in 20 mbar O2 environment. We finish the evaporation by capping in 20 mbar O2 environment for 20 minutes.
III Device assembly and wirebonding
III.1 Package
Conventional bosonic cQED architecture usually uses a monolith package, made of high-purity aluminium, to minimise seam loss [10, 55]. The chip hosting the ancillary elements is inserted into a waveguide to couple capacitively to the cavity mode and the other end of the chip is held by a clamp, which is, in turn, attached to the main package. However, this approach is not suitable for the device we use here. The on-chip ground plane and requirement for wirebonding necessitate a seam placed near the cavity wall.
Figure S5: Design of the superconducting package. (a) Top view of the main body and the bottom clamp. (b) Design of the top clamp with engineered constrictions.
Hence, our superconducting package was carefully re-designed to accommodate these requirements. Unlike conventional architecture, our package consists of three separate parts. The main body contains the cavity and the circular waveguide, where the transmon is inserted (Fig. S5(a)). This part is made of 4N6 aluminium to ensure the long lifetime of the coaxial $\lambda/4$-cavity.
The clamping structure consists of two pieces, the bottom and top clamps. As we expect the high-Q cavity mode field to be weak in the clamps volume, these parts are made of 6061 aluminium alloy to allow more precise machining and mechanical robustness.
Figure S6: Wirebonding. (a) Image of the chip on the bottom clamp after wirebonding. (b) Connection of the on-chip lines to th SMA ports. (c) Bond wires and galvanic connection between the on-chip ground plane and the aluminium bottom clamp are also realised with wirebonding.
The top clamp functions as a lid that covers the chip and shields it from the environment. Importantly, it features a 4mm-long extrusion (Fig. S5(b)), which blocks the cavity field from propagating towards SMA ports outside the heavily filtered CPW line.
The bottom clamp features a 0.5 mm-deep pocket, where the chip is located (Fig. 2(a), main text). We use wirebonding to connect the on-chip ground to the surface of the clamp surrounding it (Fig. S6(a),(c)). The package is then mounted on the MXC flange of the dilution refrigerator, effectively connecting on-chip ground plane to the common fridge ground. Similarly, we use wirebonding to galvanically connect the flux and readout line terminals to flat beryllium-copper pins, that are connected to SMA ports (Fig. S6(b)). The drive signals are supplied to the SMA ports through the coaxial lines of the dilution refrigerator.
Seam losses estimation
To avoid deterioration of the cavity lifetime, our design places the seam as far as possible from the cavity. In order to estimate cavity Q-factor due to the seam loss, we integrated the cavity field over a 0.4 mm-wide ribbon along the seam. Even though we don’t have exact data on the loss tangent of the seam, we could use that method to optimise the design based on the relative changes of the cavity Q.
Distancing the seam comes at the cost of increasing the separation between the cavity and the transmon, which could lead to decreasing their coupling strength $g$. In order to mitigate this effect, we implemented a microstrip structure to extend the transmon field and recover some of the capacitive coupling, similar to the design used in Ref. [68].
III.2 Assembly
We begin the assembly by putting a small amount of the Apiezon N cryogenic vacuum grease on the surface of the bottom clamp. We then place the chip on top and press slightly with a wooden part of the cotton swab. The grease is only present at the back part of the chip, far from the cavity.
Next, the top clamp is placed on top of the chip and fastened with 3 screws. At this point, the chip is extruding out of the clamp slightly. Finally, the cavity body and the clamping pieces are attached together with the two screws located at the bottom of the clamp.
To improve the contact at the seam in our package, we put indium wire in between the cavity and the clamping parts. The quality of this joint can be further improved with welding [51] or more advanced surface treatment [33].
Before mounting the device on the MXC flange of the dilution refrigerator, we cover all the seams between the parts with aluminium tape to ensure that the package is light-tight.
IV Calibration of transitions
The Jaynes-Cummings ladder has a rich spectrum of transitions. We calibrate the frequency and amplitude of each transition by preparing the system in one of the adjacent states and then performing spectroscopy and Rabi oscillation experiments.
For example, to calibrate the $|3-\rangle\leftrightarrow|4+\rangle$ transition the system is initialised in $|3$-$\rangle$ with the cavity and transmon on resonance. The transmon is then driven by a pulse with variable frequency and subsequently adiabatically detuned from the cavity. Finally, the transmon, now dispersively coupled to the cavity, is used to measure the parity using a standard Ramsey sequence. A change in parity indicates that drive has excited one of the four allowed transitions to states $|2\pm\rangle$ or $|4\pm\rangle$, which appear as four well-resolved features in the spectroscopy diagram. The desired $|3-\rangle\leftrightarrow|4+\rangle$ transition is identified as the one with highest frequency, in accordance with the theory.
After the transition frequency is identified, the amplitudes of excitation pulses are calibrated with power Rabi experiments. A gaussian pulse with standard deviation $\sigma$ and a total duration of $4\sigma$ is applied to the transmon line with variable amplitude. The amplitude corresponding to a $\pi$ pulse is the point in which the parity of the system is changed from odd to even. This experiment allows us to calculate the Rabi frequency $\Omega_{s}$ of the effective drive $\frac{\Omega_{s}}{2}|4+\rangle\langle 3{-}|$. The corresponding frequencies and amplitudes of each JC transition are calibrated using this method and summarised in Tab. III.
Transition $\omega_{s}/2\pi$ (GHz) $\Omega_{s}/2\pi$ (MHz) $\sigma$ (ns)
$|0\rangle\leftrightarrow|1+\rangle$ $6.8809$ $29.0$ $68$
$|0\rangle\leftrightarrow|1$-$\rangle$ $6.8566$ $25.5$ $52$
$|1-\rangle\leftrightarrow|2+\rangle$ $6.8987$ $22.3$ $44$
$|1+\rangle\leftrightarrow|2$-$\rangle$ $6.8402$ $18.9$ $48$
$|2-\rangle\leftrightarrow|3+\rangle$ $6.9074$ $24.1$ $44$
$|2+\rangle\leftrightarrow|3$-$\rangle$ $6.8315$ $19.4$ $32$
$|3-\rangle\leftrightarrow|4+\rangle$ $6.9142$ $25.9$ $44$
$|3+\rangle\leftrightarrow|4$-$\rangle$ $6.8246$ $22.0$ $24$
$|4-\rangle\leftrightarrow|5+\rangle$ $6.9201$ $25.5$ $32$
$|5+\rangle\leftrightarrow|6$-$\rangle$ $6.8143$ $22.7$
$|6-\rangle\leftrightarrow|7+\rangle$ $6.9291$ $24.5$
Table III: Calibration of the parameters of the transitions drives. The table shows the frequency of each relevant transition and its respective maximum Rabi frequency $\Omega_{s}$. $\Omega_{s}$ is defined as the coefficient of the $\frac{\Omega_{s}}{2}|\left(N+1\right)\mp\rangle\langle N\pm|$ matrix element when activated through the transmon drive. The parameter $\sigma$ is the standard deviation of the Gaussian $\pi$ pulse calibrated for each transition, with total duration of $4\sigma$. The length of the $|5+\rangle\leftrightarrow|6$-$\rangle$ and $|6-\rangle\leftrightarrow|7+\rangle$ transitions were not optimised.
We use this method to calibrate the gaussian excitation pulses used for cavity state preparation in the main text. The $\sigma$ of each pulse was swept while tracking the parity of the state after a $2\pi$ rotation. The ideal $\sigma$ is such that the parity of the final state is as close as possible to the parity of the initial state. This corresponds to a compromise between the leakage and decoherence errors that affect short and long pulses, respectively.
V Gate time and lifetime comparison
Here, we compare the resonant JC regime that we implemented to the well-known universal control techniques such as selective number-dependent arbitrary phase (SNAP) gate [32, 18], gradient ascent pulse engineering (GRAPE) [17], and echoed conditional displacement (ECD) [13] in the dispersive regime. We will make the comparison in two folds: in terms of the gate duration for generating arbitrary states on the cavity state up to Fock state $|n\rangle$ and in terms of effective lifetime of the system.
First, let us note that the dispersive regime is a special case of the JC regime where the detuning is much larger than the coupling $g$ ($\Delta=\omega_{q}-\omega_{c}\gg g$). In this regime, the main coupling term reads $\hat{H}_{\text{int}}=-\chi\hat{a}^{\dagger}\hat{a}|e\rangle\langle e|$, where $\chi\approx g^{2}/\Delta$. Notably, the dispersive regime can be further subdivided into two classes based on the ratio between $\chi$ and the fastest decoherence rate in the system, $\textnormal{max}(\Gamma_{\phi},\Gamma_{1},\kappa)$, characterising ancilla dephasing $T_{\phi}=1/\Gamma_{\phi}$ and relaxation $T_{1}=1/\Gamma_{1}$, as well as cavity lifetime $T_{1,\textnormal{cav}}=1/\kappa$. In the strong dispersive regime, the bare nonlinearity $\chi$ is considerably larger than $\textnormal{max}(\Gamma_{\phi},\Gamma_{1},\kappa)$, whereas in the weak dispersive regime $\chi\lesssim\textnormal{max}(\Gamma_{\phi},\Gamma_{1},\kappa)$. SNAP and GRAPE operate in the strong dispersive regime, where typically $\chi\sim 1\>\textnormal{MHz}$. For SNAP, by driving the qubit at a certain frequency, an arbitrary phase can be imparted on the cavity Fock state $|n\rangle$ selectively, which essentially sets the gate time as $T\sim 1/\Omega$, where $\Omega$ is the strength of the qubit drive [32]. For generating an arbitrary state within Fock state $|n\rangle$, it would involve applying the SNAP gate $n$ times, which gives $T\sim n/\Omega$ as the total gate time. However, with multi-tone qubit drives, the operation imparting arbitrary phases for all relevant Fock states can be applied simultaneously, making the gate time faster: $T_{\text{SNAP}}\sim 1/\Omega$, which is independent of $n$. And it is important to point out that for SNAP, selectivity demands $\Omega\ll\chi$, and therefore $T_{\text{SNAP}}\sim 1/\Omega\gg 1/\chi$. Another technique, GRAPE, does not require the strict selectivity of SNAP. Also, the qubit and cavity drives can be simultaneously applied while the qubit and cavity are dispersively coupled. This way, GRAPE is more general compared to SNAP, i.e., the solution space for GRAPE contains that of SNAP. This makes the gate time for GRAPE often faster than multi-tone SNAP, which is set by the dispersive coupling strength $T_{\text{GRAPE}}\sim 1/\chi\approx\Delta/g^{2}$ [47].
The technique of ECD, operating in the weak dispersive regime, enhances the nonlinearity in situ via applying strong resonant microwave drive to displace the bosonic mode far from the origin in the phase space. In this displaced frame, the dispersive four-wave-mixing interaction is transformed into three-wave-mixing, resulting in the interaction term resembling the resonant JC Hamiltonian: $\hat{H}_{\textnormal{int}}^{\textnormal{DF}}=\chi(\alpha\hat{a}^{\dagger}+\alpha^{*}\hat{a})\sigma_{z}/2$, where $\alpha$ is the displacement of the cavity in the phase space. Comparing $\hat{H}_{\textnormal{int}}^{\textnormal{DF}}$ to the Eq. (1) of the main text, one can conclude that ECD leverages the effective interaction strength $g_{\textnormal{eff}}=\chi\alpha$. This bosonic enhancement by a factor of $\alpha$, though powerful, is limited by the criticial cavity photon number [5], leading to the upper bound of the enhanced interaction of $g_{\textnormal{eff}}^{\textnormal{max}}=\sqrt{\chi E_{C}/6\hbar}$, where $E_{C}$ is the capacitive energy of the transmon, approximately equal to the transmon anharmonicity. For the typical transmon designs, $E_{C}/2\pi\hbar$ is in the range of $\textnormal{200}-\textnormal{300 MHz}$ in order to suppress the charge noise. In Ref. [13], where $\chi/2\pi=\textnormal{33 kHz}$ and $E_{C}/2\pi\hbar=\textnormal{193 MHz}$, $g_{\textnormal{eff}}^{\textnormal{max}}\approx\textnormal{1 MHz}$, which sets the lower bound for the gate time of one ECD gate to be $t_{\textnormal{ECD}}\geq 1/g_{\textnormal{eff}}^{\textnormal{max}}$. Furthermore, universal control in this regime involves sequential applications of ECDs and qubit rotations, where the complexity of applying an arbitrary gate of dimension $n$ scales with the number of applications $n$.
On the other hand, protocol for generating arbitrary superpositions of Fock states has also been realised utilizing sequential off- and on-resonant (JC regime) qubit-cavity regimes [38, 20]. Here, each JC ladder transition $|e,n-1\rangle\leftrightarrow|g,n\rangle$ has a Rabi rate $\Omega\sqrt{n}$, making each step $t_{n}\sim 1/(g\sqrt{n})$ long. For an operation involving $n$ ladders, the total gate time is approximately $T_{\text{JCS}}=\sum_{n}t_{n}\sim\sqrt{n}/g$. Importantly, this protocol was further improved by incorporating qubit drives with multi-tone frequencies, which drive all relevant transitions at the same time [63], ultimately resulting in $T_{\text{JC}}=\max_{n}t_{n}\sim 1/g$. If we compare this to the dispersive regime (GRAPE), we have $T_{\text{JC}}\sim(g/\Delta)\times T_{\text{GRAPE}}$. Considering that $g/\Delta\approx 0.05-0.2$, the gate time utilizing JC regime is 5205-20 times faster.
For effective lifetime comparison, let us assume that the major contribution comes from qubit energy decay and dephasing, where the corresponding lifetimes are denoted as $T_{1}$ and $T_{\phi}$, since the cavity lifetime is normally much longer. For SNAP, GRAPE, and ECD, the qubit is utilised differently during a particular protocol. However, the error probabilities due to the qubit can be approximated in the same way. In particular, the error probability related to qubit decay is roughly estimated to be $\epsilon_{1}\sim p_{e}\times T/T_{1}$, where $p_{e}$ is the average probability of the qubit being in the excited state (when the decay happens) during the protocol. This approximation also applies for dephasing $\epsilon_{2}\sim 2\sqrt{p_{e}p_{g}}\times T/T_{\phi}$, which is affecting the qubit state in superposition, the maximum of which happens for equal superposition $p_{g}=p_{e}=0.5$, leading to maximum error $\sim T/T_{\phi}$. Therefore, the total error goes as $\epsilon\sim T(p_{e}/T_{1}+2\sqrt{p_{e}p_{g}}/T_{\phi})$. One can define an effective lifetime for the system $T_{\text{sys}}$ in relation to the error and total gate time as $\epsilon\sim T/T_{\text{sys}}$ such that $T_{\text{sys}}\sim(p_{e}/T_{1}+2\sqrt{p_{e}p_{g}}/T_{\phi})^{-1}$.
In the case of the JC regime, as written in the main text the effective decay lifetime of the coupled system is approximately twice that of the qubit $2T_{1}$, while the dephasing (depolarisation) is $T_{\phi}$, as the JC eigenstates contain equal superposition of the qubit’s ground and excited states. Consequently, this results in effective lifetime $T_{\text{sys}}\sim(1/(2T_{1})+1/T_{\phi})^{-1}$ for the system. A comparison for the lifetime is generally hard to make since it is protocol specific. However, one can see that for average qubit probability $p_{e}\sim 0.5$ in the dispersive regime, the system lifetime is equal to the JC regime.
An important point worth stressing is that we detune the qubit after the protocol such that the qubit is decoupled from the cavity, and therefore the cavity’s lifetime is back to its original value. Even though the resonant control protocol introduces the small overhead due to the time required to adiabatically detune the system from resonance, this overhead is constant (does not scale with the gate complexity) and reasonably short even on the timescale of the current system lifetimes ($\sim\textnormal{200}$ ns).
The resonant scheme achieves fast gate time without compromising the lifetime of the system. Therefore, we argue that the JC protocol that we implement here offers the most favourable strategy for universal control of a bosonic mode.
Our method can be further improved by exploring the deep-strong (DS) and ultra-strong (US) coupling regimes of the JC Hamiltonian, providing access to significantly higher values of coupling strength $g$. However, the rotating wave approximation of the quantum Rabi model [49], used in current analysis (see Eq. (1) of the main text), will not be applicable anymore in those regimes. Therefore, new strategies are to be developed for universal control of JC Hamiltonian in DS and US regimes.
Other methods, such as tailored nonlinear Hamiltonians beyond JC dynamics, can also be explored. For instance, the cubic interactions [14] have been demonstrated on a low-Q oscillator by directly terminating the planar cavity with a nonlinear element. The strong nonlinearity inherent to such bosonic mode allows operations on the timescale of tens of nanoseconds. However, such an element is not an effective quantum memory, as the nonlinearity cannot be dynamically decoupled from the cavity, leading to always-on high decoherence rate. Furthermore, this scheme can only operate in the bounded area of the phase space ($|\alpha|<1.5$), outside of which the control is undermined by the static Kerr nonlinearity. Therefore, at this stage such an approach still requires further investigation in order to be a general-purpose tool to achieve universal control of a bosonic mode.
VI Preparation of arbitrary bosonic states
We used the resonant control protocol to demonstrate the preparation of arbitrary states in the cavity with an analytical control sequence. The sequence consisted of activating the sidebands sequentially using variable-amplitude, Gaussian-shaped pulses with $4\sigma$ duration, with the parameters given in the Tab. III. This method allows for straightforward preparation of any Fock state and their superpositions in the cavity, as shown in Fig. S7.
Figure S7: Qudit arbitrary state preparation. Wigner plots showing diverse qudit states prepared by sequentially exciting sideband transitions in the Jaynes-Cummings ladder.
Due to the lack of appropriate quantum amplifier at the specific moment of this experiment, we don’t have single-shot readout of the transmon. Thus, the parity signal was calibrated by comparing it to a $1\textnormal{D}$ Wigner measurement of cavity vacuum $|0\rangle$, which is independently measured beforehand. The Wigner plots closely resemble the target states in shape, though with a reduction in contrast. This indicates that the state preparation is mostly limited by the transmon decoherence.
VII Arbitrary gate on logical states of binomial encoding
As shown in Sec. VI, resonant control protocol enables arbitrary state preparation. Here, we show that an arbitrary gate on the logical states of binomial encoding of the lowest order [50] can be achieved within the framework of the resonant control protocol. This follows as Givens rotations can be utilized to realise $\hat{R}_{x}^{\textnormal{L}}(\theta)$, $\hat{R}_{y}^{\textnormal{L}}(\theta)$, and $\hat{R}_{z}^{\textnormal{L}}(\theta)$ on the subspace $\big\{|\mathord{\downarrow}\rangle^{\textnormal{L}}=|2\rangle,|\mathord{\uparrow}\rangle^{\textnormal{L}}=(|0\rangle+|4\rangle)/\sqrt{2}\big\}$. The rotations can be written explicitly in terms of Givens rotations as
$\hat{R}_{x}^{\textnormal{L}}(\theta)=\hat{\mathcal{R}}_{0,2}(\zeta_{1})\hat{\mathcal{R}}_{2,4}(\zeta_{2})\hat{\mathcal{R}}_{0,2}(\zeta_{1}),$ (S7a)
$\hat{R}_{y}^{\textnormal{L}}(\theta)=\hat{\mathcal{R}}^{-\pi/2}_{0,2}(\zeta_{1})\hat{\mathcal{R}}^{\pi/2}_{2,4}(\zeta_{2})\hat{\mathcal{R}}^{-\pi/2}_{0,2}(\zeta_{1}),$ (S7b)
$\hat{R}_{z}^{\textnormal{L}}(\theta)=\hat{\mathcal{R}}_{0,4}(-2\theta),$ (S7c)
where the parameters $\zeta_{1}=2\arctan\big(\tan(\theta/4)/\sqrt{2}\big)$, $\zeta_{2}=4\arcsin\big(\sin(\theta/4)/\sqrt{2}\big)$, and the notation $\hat{\mathcal{R}}^{\gamma}_{n_{A},n_{B}}(\alpha)$ means Givens rotation between the Fock states $|n_{A}\rangle$ and $|n_{B}\rangle$ by an angle $\alpha$ as described in the main text, where the pulse is phase-shifted by $\gamma$.
Thus, arbitrary control on the logical states of the binomial encoding is enabled via trivial combinations of the Givens rotations.
VIII Parametric modulation
Although not required for the resonant control protocol, our device architecture is also capable of providing broadband parametric flux modulation of the SQUID.
We show this feature experimentally by using flux pumps to activate parametric resonances between the cavity and transmon with tunable strength. Sweeping the pump frequency $\omega_{m}$ around the transmon-cavity detuning of $\Delta\approx 85\,\textnormal{MHz}$ for a variable duration results in the continuous exchange of the transmon excitation with the cavity, as shown by the chevron pattern in Fig. S8(a). The resonant oscillations have a period of $\pi/g_{\textnormal{eff}}$, where $g_{\textnormal{eff}}$ is the parametric coupling strength.
Figure S8: Protected flux-parametric modulation. (a) Vacuum Rabi oscillations activated by a flux sideband. The transmon-cavity detuning is $\Delta=\textnormal{85 MHz}$. After being excited with a $\pi$ pulse, a flux pump with $\omega_{m}/2\pi\approx\Delta$ creates a sideband that activates the photon exchange between transmon and cavity for a variable interaction time. (b) Resonant vacuum Rabi oscillations as a function of the flux pump amplitude. The speed of the Rabi oscillations can be continuously tuned in the $g_{\textnormal{eff}}/2\pi=\textnormal{0}\div\textnormal{6.57 MHz}$ range. (c) Transmon spectroscopy for the variable values of $\Delta$ under a flux pump $\varphi_{e}\propto\cos\big(2\pi\Delta t\big)$. The flux pump activates the parametric resonance between the transmon and the cavity, as seen from the frequency split that appears on every sideband due to the avoided crossing.
The parametric coupling strength $g_{\textnormal{eff}}$ can be tuned on demand by changing the flux pump amplitude. Fig. S8(b) shows that the resonant oscillations become faster with larger flux pump amplitude (expressed in units of current), corresponding to an increase in coupling strength. The pump amplitude $I_{p}$ is estimated from the room-temperature input current after going through $26\,\textnormal{dB}$ attenuation and treating the flux line as a short. Fitting the oscillations to a cosine function for each pump amplitude indicates a variable $0\leq g_{\textnormal{eff}}\leq 6.57\,\textnormal{MHz}$. However, increasing the pump amplitude beyond the maximum coupling point leads to lower $g_{\textnormal{eff}}$. This is consistent with the expected analytical trend $g_{\textnormal{eff}}=g\times J_{1}(I_{p}/\nu)$, where $g$ is the static coupling, $J_{1}$ is a Bessel function of the first kind and $\nu$ is proportionality factor [12].
[…]
$\nu=\left(\frac{2\pi}{\Phi_{0}}\frac{d\omega_{T}}{d\varphi_{e}}\frac{L}{\omega_{m}}\right)^{-1},$ (S8)
where $\omega_{T}/2\pi=6.783\,\textnormal{GHz}$ is the transmon parking frequency and $L$ is the effective mutual inductance with the SQUID. From the fitting value of $\nu=53\,\mu\textnormal{A}$ we can calculate $L=0.211\,\Phi_{0}/\textnormal{mA}$, which is considerably lower than the DC bias mutual inductance of $L_{DC}=0.387\,\Phi_{0}/\textnormal{mA}$. This lower tunability can be partially attributed to the flux pump not fully driving the differential mode of the SQUID loop [43]. In further device iterations, the mismatch can be mitigated by incorporating a common-mode rejection filter on the flux line [40].
We test strong parametric resonances with flux pumps at higher modulation frequency. This is shown with transmon spectroscopy experiments at different parking detunings $\Delta=63~\textnormal{to}~232\,\textnormal{MHz}$ under a continuous flux pump with frequency $\omega_{m}=\Delta$. The pump creates flux sidebands at frequencies $\bar{\omega}_{T}+N\omega_{m}$, where the average transmon frequency $\bar{\omega}_{T}$ approximately coincides with the parking frequency $\omega_{T}$ for low pump amplitudes. In all spectroscopy plots (Fig. S8(c)), the rightmost feature corresponds to the $N=1$ sideband, which coincides with the cavity frequency and presents a frequency split corresponding to their avoided crossing. The next feature to the left corresponds to the transmon average frequency ($N=0$), and the lower N=1N=-1 and N=2N=-2 sidebands can be seen within range of the spectroscopy for some DC biases. All features display the same frequency splitting indicating the strong coupling of the $N=1$ sideband to the cavity for pump frequencies as high as $232\,\textnormal{MHz}$.
IX Quantum state reconstruction
Here, we explain the procedure for reconstructing a quantum state, represented by its density matrix, with truncation dimension $D$, given the Wigner measurement data. An arbitrary $D$ dimensional quantum state $\rho$ is fully characterised by its $D^{2}-1$ independent real parameters, which we write in a vector form $\vec{Y}$. There are many choices for the parametrisation of $\rho$. We take the first D1D-1 diagonal elements of $\rho$ as well as the $D^{2}-D$ real and imaginary off-diagonal (upper triangular) elements. This way, the vector $\vec{Y}$ consists of real elements, which fully characterise the density matrix $\rho$.
On the other hand, the Wigner measurement that we performed is expressed as $W(\alpha)=\text{tr}(\hat{P}\hat{D}(\alpha)^{\dagger}\rho\hat{D}(\alpha))$ with $\hat{P}=e^{i\hat{a}^{\dagger}\hat{a}\pi}$ being the parity operator and grid-based displacement points $\{\alpha\}$. Let us express all the Wigner data in a vector form $\vec{X}$. Consequently, the input state and Wigner data are related linearly, which can be written in a single matrix equation $\vec{X}=M\vec{Y}+\vec{V}$, where $M$ and $\vec{V}$ denote a mapping matrix and a constant vector, which can easily be computed given the choice of displacements $\{\alpha\}$.
The state reconstruction is carried out in a few steps. First, we perform linear inversion (LI) to get the estimated state parameters $\vec{Y}_{\text{est}}$, given the Wigner measurement data $\vec{X}$: $\vec{Y}_{\text{est}}=M^{+}(\vec{X}-\vec{V})$, where $M^{+}=(M^{\dagger}M)^{-1}M^{\dagger}$ is the left Moore-Penrose pseudoinverse of the matrix $M$. Then, we construct the density matrix $\rho_{\text{LI}}$ based on the estimated parameters. As errors are present in experiments, this density matrix might not be physical, i.e., it might contain negative eigenvalues. Thus, given $\rho_{\text{LI}}$, we employ Bayesian inference (details below) to obtain a posterior probability distribution that accurately represents experimental data. From this distribution, we can sample density matrices and compute the mean fidelity and its standard deviation.
IX.1 Bayesian inference
A popular technique to obtain a physical density matrix from experimental data in the presence of noise is maximum likelihood (ML), which gives us the closest physical density matrix $\rho_{\text{ML}}$ from the obtained density matrix $\rho_{\text{LI}}$ via linear inversion [62]. However, it is statistically less accurate. For example, it often leads to $\rho_{\text{ML}}$ having zero eigenvalues [6, 30]. This is an issue as zero eigenvalues are interpreted as zero probabilities if we were to measure in the corresponding basis, which is not a valid conclusion as we perform finite measurement repetitions.
To address this issue, we use Bayesian inference, which is a statistically more accurate approach. Through Bayes’ rule, it gives us a posterior distribution, explicitly showing uncertainty, given the measured experimental data ($\rho_{\text{LI}}$). The posterior distribution allows us to sample any functions of $\rho$ (including itself). For exact implementation of Bayesian inference, we follow an efficient protocol proposed in Ref. [46]. When implementing the protocol, we use the following parameters: $\alpha=1$, which establishes a uniform prior as we do not assume initial preference for the density matrix; the variance of the pseudo-likelihood $\sigma=1/N$ with $N=\#\>\cdot\>(D^{2}-1)$, where #\# is the number of experimental repetition; we also generate $2^{10}$ Markov Chain Monte Carlo samples and apply a thinning parameter of $2^{7}$ to mitigate correlations within the sample chain.
For example, we use this procedure to sample a large number of possible physical density matrices $\{\rho_{\text{BI},i}\}$ and compute their fidelities $F_{i}=(\text{tr}(\sqrt{\sqrt{\rho_{\text{BI},i}}\>\rho_{\text{tar}}\>\sqrt{\rho_{\text{BI},i}}}))^{2}$ to a target state $\rho_{\text{tar}}$. The fidelity and its error presented in the main text are the average fidelity and its standard deviation.
The quantum state reconstruction and Bayesian inference routines used in that work are available on GitHub, github.com/Qcrew/Valadares-Dorogov-FAR, along with the raw experimental data (averaged).
X Protocol simulation and error budget
We estimate the effect of different error sources over the resonant control protocol by solving the master equation of the system.
First, we simulate the state preparation of states $|\psi_{1}\rangle=\frac{1}{\sqrt{2}}\left(|0\rangle+|3\rangle\right)$ and $|\psi_{2}\rangle=\tfrac{1}{2}|0\rangle+\tfrac{\sqrt{3}}{\sqrt{8}}\left(i|2\rangle+|4\rangle\right)$ using sequential gaussian pulses for each transition. The system Hamiltonian was reconstructed using the experimentally calibrated transitions from Tab. III. The cavity lifetime $T_{1,\textnormal{cav}}$, transmon lifetime $T_{1}$ and pure dephasing $T_{\phi}=\left(\frac{1}{T_{2}}-\frac{1}{2T_{1}}\right)^{-1}$ are implemented with the respective jump operators. The adiabatic decoupling is taken as ideal and implemented with a change of basis operation. The Wigner tomography executed in the dispersive regime is also reproduced in simulation with experimental parameters, namely a dispersive shift $\chi/2\pi=1.62\,\textnormal{MHz}$, a parity-mapping time of $272\,\textnormal{ns}$ and gaussian $\pi/2$ pulses with $\sigma_{\pi/2}$ of $8\,\textnormal{ns}$ and total duration of $4\sigma$. The simulated tomography is then rescaled following the amplitude of the Wigner function of vacuum obtained with consistent simulation parameters. The state fidelity is then estimated from density matrix reconstruction following the procedure described in Sec. IX.
Figure S9: Simulated state preparation fidelities. Simulated state fidelity of $|\psi_{1}\rangle=\frac{1}{\sqrt{3}}\left(|0\rangle+|3\rangle\right)$ as a function of transmon $T_{1}$ and $T_{\phi}$.
The state preparation is simulated for transmon $T_{1}=10\,\mu\textnormal{s}$ and $15\,\mu\textnormal{s}$, and $T_{2}$ in the $\left[1.3,2.8\right]\,\mu\textnormal{s}$ range. The cavity lifetime is fixed at $T_{1,\textnormal{cav}}=500\,\mu\textnormal{s}$. We first identify that the transmon is highly sensitive to changes in dephasing, as seen from the reconstructed fidelity as function of $T_{\phi}$ in Fig. S9: the fidelity of $|\psi_{1}\rangle$ varies in a large range $87.2\,\%,93.8\,\%]$ ($87.3\,\%,93.8\,\%]$) for $T_{1}=10\,\mu\textnormal{s}$ ($15\,\mu\textnormal{s}$). These values also show the protocol is weakly sensitive to variations in transmon $T_{1}$ in this parameter range. Similar conclusions about $T_{1}$ and $T_{2}$ sensitivity can be drawn for state $|\psi_{2}\rangle$. Simulations with $T_{2}=2.7\,\mu\textnormal{s}$ and $T_{1}=15\,\mu\textnormal{s}$ ($T_{\phi}\approx 2.97\,\mu\textnormal{s}$) lead to $93.2\,\%$ and $87.7\,\%$ for states $|\psi_{1}\rangle$ and $|\psi_{2}\rangle$, respectively, reproducing well the experimental fidelities of $93\pm 3\,\%$ and $89\pm 2\,\%$.
Included error sources $|\psi_{1}\rangle$ fid. (%) $|\psi_{2}\rangle$ fid. (%)
Finite pulse length $98.4$ $95.1$
+ $T_{1}$ $98.3$ $94.4$
+ $T_{\phi}$ $93.2$ $87.7$
Table IV: Error budget for state preparation. State fidelity for $|\psi_{1}\rangle=\frac{1}{\sqrt{2}}\left(|0\rangle+|3\rangle\right)$ and $|\psi_{2}\rangle=\tfrac{1}{2}|0\rangle+\tfrac{\sqrt{3}}{\sqrt{8}}\left(i|2\rangle+|4\rangle\right)$ states with progressive addition of error sources.
We calculate the fidelities of $|\psi_{1}\rangle$ and $|\psi_{2}\rangle$ while progressively including more sources of error in the simulation (Tab. IV). The results indicate that the dephasing $T_{\phi}$ has the largest impact on state preparation fidelity, followed by coherent errors due to finite pulse length, and transmon $T_{1}$, in that order. From these simulations, we expect that the quality of state preparation can be significantly enhanced even with modest improvements in the transmon $T_{\phi}$.
We simulate the Givens rotations between Fock states $|1\rangle$ and $|4\rangle$ using the same procedure, with the addition of the optimised $|1+\rangle\leftrightarrow|3$-$\rangle$ state transfer drive. The simulated cavity photon distributions are compared with the experimental reconstructed density matrix in Fig. 4c in the main text, showing a close match.
Simulations indicate that the experimental state fidelities after Givens rotations are significantly affected by coherent errors. When neglecting the transmon $T_{1}$ and $T_{2}$, the state transfer $|1+\rangle\to|3$-$\rangle$ shows a fidelity of $95.0\,\%$ when the system is initialised in $|1+\rangle$. For the same drive, the returning $|3-\rangle\to|1+\rangle$ path shows a fidelity of $85.8\,\%$ when the system is initialised in $|3$-$\rangle$. We attribute these imperfections to the mismatch between the ideal Jaynes-Cummings model used in the drive optimization and the experimental Hamiltonian parameters, which leads to a difference on the order of $\approx 1\,\textnormal{MHz}$ in the frequencies of higher transitions. We expect that further improvements in our calibration and optimization methods will render significant improvement in the Givens rotations fidelities.